# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

### v2

@@ 1 Introduction
In the following, we will start by briefly introducing the mathematical formulation of PCA-clustering (more details can be found in [16, 17, 18]), followed by demonstrating that the main bottleneck of PCA-clustering is successfully mitigated without increasing the leading order of the cost scaling, by the two modifications of its mathematical formulation: (i) by upgrading the original manifold distance loss from PCA (being a semi-norm) to the weighted combination of manifold distance and squared Euclidean distances, hereby making the clustering loss function to a weighted norm (metrisation step); (ii) and by including the Shannon entropy regularization on the cluster affiliation probability measures, making the resulting clusterings entropy-optimal (entropy regularization step). Finally, the resulting unsupervised Entropy Optimal Clustering algorithm (EOMC) will be investigated mathematically, and compared to the most popular linear and nonlinear manifold learning algorithms on synthetic benchmark examples, and on the outputs of the Lorenz-96 model in chaotic and very chaotic regimes.
@@ 2.1 Mathematical formulation of the PCA-clustering
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\tilde{\mathcal{L}}$
$\tilde{\mathcal{L}}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\|X(:,t)-\sum_{k=1}^{K}\gamma(k,t)X^{\textrm{rec},(k)}(:,t)\|_{2}^{2},$
@@
Substituting (1) in (2.1) and applying the Jensen inequality, we obtain that the values of $\gamma$ and $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ can be approximated by minimizing the Jensen’s upper bound $\mathcal{L},\mathcal{L}\geq\tilde{\mathcal{L}}$ of the problem (2.1):
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}$
$\mathcal{L}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2},$
@@
PCA-clustering algorithm iteratively finds locally-optimal solutions of the problem (2.1) using two of its mathematical properties: (i) for a fixed $\gamma$, this problem has an analytic solution with respect to manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$, provided by the cluster weighted means and the dominant eigenvectors of the cluster-weighted covariance matrices; and (ii) for the fixed manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and for each of the $t$, $\gamma(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2}$, and $\gamma(k,t)=0$ for all $k\neq k^{*}$ delivers the analytic $\gamma$-solutions of (2.1) [^1] Please note that since the optimal $\gamma$ in this solution takes only $0/1$-values, Jensen inequality becomes the Jensen equality and, hence, $\mathcal{L}=\tilde{\mathcal{L}}$. PCA-clustering starts with a random intialization of $\gamma$ and iteratively repeats these analytic steps (i) and (ii), resulting in the monotonic convergence of $\mathcal{L}$, with overall iteration complexity scaling linearly in $T$ [16, 17, 18].
However, the Lemmas 1 and 2 below gives rise to a significant mathematical difficulty, when applying PCA-clustering to approximation of low-dimensional nonlinear manifolds.
@@
$\mathrm{dim}(\mathrm{Ker}(B))=D-\mathrm{rank}(P)=D-d$
@@
From Lemmas 1 and 2 follows the first major problem of PCA-clustering: according to the Lemmas, the kernel of the projection operator is non-empty when $d<D$ (Lemma 1), and any perturbations inside of the kernel do not have any effect on the loss (Lemma 2). Hence, loss function in (2.1) is a semi-norm, not penalizing errors in the directions that are orthogonal, or close to orthogonal with respect to the manifolds $\mathcal{T}_{k}$. As will be demonstrated on examples below (see, e.g., Fig. 2C and the left panel of Fig. 3C), this can lead to the PCA-clustering minimizers $\mathcal{T}_{k}$ that become orthogonal, or close to orthogonal, with respect to the actual nonlinear low-dimensional manifold - hereby ignoring the actual manifold data, and achieving a small value of $\mathcal{L}$ at the same time. The second problem of PCA-clustering is induced by the exact $0/1$ minimizer in the step (ii) described above: it confines the resulting manifold reconstruction $X^{\textrm{rec}}(:,t)$ in (2) to piece-wise linear functions only, not allowing to achieve smooth convex interpolations of nonlinear manifolds that should be possible with the original optimization problem formulation (2.1). But, this original formulation (2.1) does not allow obtaining the cheaply-computable analytic linearly-scalable solutions (i)-(ii) that are used in the PCA-clustering algorithm. And, solving the original problem (2.1) - being highly-nonlinear, subject to multivariate and polynomial constraints - with standard numerical tools from optimization theory would require very expensive numerical algorithms that have the exponential worst-case scaling [19].
@@ 2.2 Entropy-Optimal Manifold Clustering (EOMC)
To mitigate these two central problems of PCA-clustering, the two following modifications of the loss function $\mathcal{L}$ in (2.1) will be adopted:
  • Loss metrisation: to fix the problem identified by the Lemmas 1 and 2, we will introduce an additional scalar hyper-parameter $\alpha\geq 0$ and add an addtional term $\alpha\|X(:,t)-\mu_{k}\|_{2}^{2}$ to the loss of each of the $t$ instances for all $k=1,\dots,K$. As will be proven below, this modification upgrades the semi-norm loss of (2.1) to the $\alpha$-weighted norm, fixing the problem with the data points in the directions that are close to orthogonal to $\mathcal{T}_{k}$.
  • Entropy regularization: to avoid the problem with a piece-wise linearity restriction - and to allow for the $\gamma$-solutions to be not only $0/1$ (and, to achieve this without returning back to the original formulation (2.1) that would result in exponentialy-scaling worst case numerics) - we can use the fact that the columns of $\gamma$ can be interpreted as $t$-dependent $K$-dimensional probability measures. We can add the normalized negative Shannon entropy regularization terms $\beta\gamma(k,t)\textrm{log}_{K}\left(\gamma(k,t)\right)$ (with a scalar hyper-parameter $\beta\geq 0$) to the loss, to optimally-adjust the entropies of these probability measures during learning [^2] Base $K$ of the logarithm is used to normilize the entropy values to the interval $\left[0,1\right]$. This modification is motivated by the entropic learning methods that allow training the entropy-optimal probability measures by tuning this hyper-parameter $\beta$ in the loss. The $0/1$ $\gamma$-solution of the PCA-clustering is a minimum entropy measure, and is attained when setting $\beta=0$ [20, 21, 22, 23, 24].
@@
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}^{\textrm{EOMC}}$
$\mathcal{L}^{\textrm{EOMC}}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right],$
@@
Please note that the PCA-clustering [16, 17, 18] is a special case of (2.2), attained when setting $\alpha=0$ and $\beta=0$. Standard linear PCA is also a special case of (2.2), when selecting $K=1,\alpha=0,\beta=0$.
@@
$x^{\dagger}A^{\alpha}x=x^{\dagger}\left((1+\alpha)I_{D}-P\right)x$
$x^{\dagger}A^{\alpha}x=(1+\alpha)x^{\dagger}I_{D}x-x^{\dagger}Px$
$x^{\dagger}A^{\alpha}x=(1+\alpha)\|x\|^{2}-\|Px\|^{2}$
@@
$x^{\dagger}A^{\alpha}x=\|x\|^{2}+\alpha\|x\|^{2}-\|Px\|^{2}$
$x^{\dagger}A^{\alpha}x=(\|x\|^{2}-\|Px\|^{2})+\alpha\|x\|^{2}$
@@
Finally, as proven in the Theorem 1 below, optimization problem (2.2) can be solved through a sequence of analytically-solvable steps - and without exceeding the computational cost of the original PCA-clustering algorithm for (2.1), i.e., with the overall leading iteration cost scaling of $\mathcal{O}(T)$.
@@
Let $X\in\mathbb{R}^{D,T}$, and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$ are fixed. Then, the EOMC problem (2.2) has the following properties:
@@
    Step 1 of EOMC-algorithm, analytic solutions for $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$: if $\sum_{t=1}^{T}\gamma(k,t)>0$, then for a fixed $\gamma$ that satisfies the constraints in (2.2), the solution of (2.2) is provided by:
@@
    where $\underset{d}{\overline{\mathrm{eigvec}}}\left(A\right)$ denotes an operation of computing $d$ dominant eigenvectors of A (i.e., the eigenvectors that correspond to the $d$ largest eigenvalues of the symmetric positive-semidefinite operator $A$) and putting them in the $D\times d$ orthogonal matrix column-wise. Operator $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ is defined as $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)=\frac{\sum_{t=1}^{T}\gamma(k,t)\left(X(:,t)-\mu_{k}^{*}\right)\left(X(:,t)-\mu_{k}^{*}\right)^{\dagger}}{\sum_{t=1}^{T}\gamma(k,t)}$. Cost of (8-9) computation scales as $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$.
@@
    Step 2 of EOMC-algorithm, analytic solutions for $\gamma(k,t)$: let $g(t)=\left(g_{1}(t),\dots,g_{K}(t)\right)$, where $g_{k}(t)=\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)$, and let $g_{\min}(t)$ be the infinum of $g(t)$ for a fixed $t$. Then, for fixed $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and when $\beta>0$, the solution of (2.2) is given by:
@@
    If $\beta=0$, then the solution of (2.2) is given by $\gamma^{*}(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\left(g_{k}(t)-g_{\min}(t)\right)$, and $\gamma^{*}(k,t)=0$ for all $k\neq k^{*}$. Cost of computing $\gamma^{*}$ with (8-9) scales as $\mathcal{O}\left(TK\left(D+1\right)\right)$.
@@
    Monotonicity and convergence of EOMC-algorithm: starting with a randomly-generated $\gamma$ that satisfies constraints in (2.2), and iteratively repeating the above Step 1 and Step 2, results in monotonic decrease of the function $\mathcal{L}^{\textrm{EOMC}}$, converging to a local optimum of (2.2). The overall computational iteration cost of EOMC-algorithm scales as $\mathcal{O}\left(TK\left(D+1\right)^{2}+KdD^{2}\right)$
@@
    For fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$, for every $\mu_{k}$, $k=1,\dots,K$, according to the Lemma1 the problem (2.2) is a strictly-convex unconstrained problem with a unique minimum
    $\sum_{t=1}^{T}\gamma(t,k)\left(X(:,t)-\mu_{k}^{*}\right)=0,$
    that after re-arranging of terms provides (8) if $\sum_{t=1}^{T}\gamma(k,t)>0$. If $\sum_{t=1}^{T}\gamma(k,t)=0$ then any value of $\mu_{k}^{*}$ is a solution. Next, we fix these $\mu_{k}^{*}$, for all $k=1,\dots,K$, and consider the solution of (2.2) with respect to $\mathcal{T}_{k}$, for fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$. Defining the $d\times d$ matrix of Lagrange multipliers $\Lambda$, we can rewrite (2.2) in the Euler-Lagrange form
@@
    $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}$ $=$ $I_{d}.$
    Substituting (13) into (12), taking trace and deploying algebraic transformations that make use of the trace operation properties, we obtain that $\Lambda$ in (13) is a diagonal matrix of dominant eigenvalues of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (trace of $\Lambda$ should maximize the trace of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ after projection on the orthogonal subspace $\mathcal{T}_{k}$). Hence, $\mathcal{T}_{k}$ corresponds column-wise to the $d$ dominant eigenvectors of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ and $\mathcal{T}_{k}^{*}=\underset{d}{\overline{\mathrm{eigvec}}}\left(\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)\right)$. Summing up the cost of computing the weighted means and weighted covariances for each cluster with the cost of computing the $d$-dominant eigenvectors of symmetric positive-semidefinite matrix $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (that is less then a cost of full diagonalization, and allows applying iterative Krylov methods with cost scaling of $\mathcal{O}\left(KdD^{2}\right)$), we obtain the cost scaling of $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$ for Step 1 .
@@
    These properties follow from applying Lemma 2.1 and Lemma 2.6 in [24] to (2.2). The cost of computing Step 2 consists of $\mathcal{O}\left(TKD\right)$ (computing $K$ elements of vector $g(t)$ for all $t$), and adding the costs of $\mathcal{O}\left(TK\right)$ computations of (10) for every $t$.
@@
    As follows from the above proofs for items (1.) and (2.), each of the Step 1 and Step 2 leads to a monotonic decrease of the function value $\mathcal{L}^{\textrm{EOMC}}$ in (2.2). It is straightforward to validate that this sequence is bounded from below with $\mathcal{L}^{\textrm{EOMC}}\geq-\beta$. Hence, since iterative repetition of these steps also generates a monotonically-decreasing sequence - and it is bounded from below with $\beta$ in $\mathcal{R}^{1}$ - this monotonically-decreasing sequence is converging to some finite $\mathcal{L}^{\textrm{EOMC}}$ if $\beta<\infty$, s.t. $\|\mathcal{L}^{\textrm{EOMC}}\|<\infty$.

### v3

@@ 1 Introduction
In the following, we will start by briefly introducing the mathematical formulation of PCA-clustering (more details can be found in [16, 17, 18]), followed by demonstrating that the main bottleneck of PCA-clustering is successfully mitigated without increasing the leading order of the cost scaling, by the two modifications of its mathematical formulation: (i) by upgrading the original manifold distance loss from PCA (being a semi-norm) to the weighted combination of manifold distance and squared Euclidean distances, hereby making the clustering loss function to a weighted norm (metrisation step); (ii) and by including the Shannon entropy regularization on the cluster affiliation probability measures, making the resulting clusterings entropy-optimal (entropy regularization step). Next, the resulting unsupervised Entropy Optimal Clustering algorithm (EOMC) will be investigated mathematically, and compared to the most popular linear and nonlinear manifold learning algorithms on synthetic benchmark examples, and on the outputs of the Lorenz-96 model of simplified atmospheric dynamics in chaotic and very chaotic regimes. Finally, the findings from Lorenz-96 model analysis will be validated applying EOMC to a much more advanced modified Hasegawa-Wakatani (mHW) plasma turbulence model from Magnetohydrodynamics (MHD).
@@ 2.1 Mathematical formulation of the PCA-clustering
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\tilde{\mathcal{L}}$ (4)
$\tilde{\mathcal{L}}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\|X(:,t)-\sum_{k=1}^{K}\gamma(k,t)X^{\textrm{rec},(k)}(:,t)\|_{2}^{2},$ (5)
@@
Substituting (1) in (5) and applying the Jensen inequality, we obtain that the values of $\gamma$ and $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ can be approximated by minimizing the Jensen’s upper bound $\mathcal{L},\mathcal{L}\geq\tilde{\mathcal{L}}$ of the problem (4-7):
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}$ (8)
$\mathcal{L}$ $=$ $\frac{1}{T}\sum_{k=1}^{K}\sum_{t=1}^{T}\gamma(k,t)\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2},$ (9)
@@
PCA-clustering algorithm iteratively finds locally-optimal solutions of the problem (8-11) using two of its mathematical properties: (i) for a fixed $\gamma$, this problem has an analytic solution with respect to manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$, provided by the cluster weighted means and the dominant eigenvectors of the cluster-weighted covariance matrices; and (ii) for the fixed manifold parameters $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and for each of the $t$, $\gamma(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\|\left(X(:,t)-\mu_{k}\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\left(X(:,t)-\mu_{k}\right)\|_{2}^{2}$, and $\gamma(k,t)=0$ for all $k\neq k^{*}$ delivers the analytic $\gamma$-solutions of (8-11) [^1] Please note that since the optimal $\gamma$ in this solution takes only $0/1$-values, Jensen inequality becomes the Jensen equality and, hence, $\mathcal{L}=\tilde{\mathcal{L}}$. PCA-clustering starts with a random intialization of $\gamma$ and iteratively repeats these analytic steps (i) and (ii), resulting in the monotonic convergence of $\mathcal{L}$, with overall iteration complexity scaling linearly in $T$ [16, 17, 18].
However, the Lemmas 1 and 2 below give rise to a significant mathematical difficulty, when applying PCA-clustering to approximation of low-dimensional nonlinear manifolds.
@@
$\mathrm{dim}(\mathrm{Ker}(B))=D-\mathrm{rank}(P)=D-d.$ (15)
@@
From Lemmas 1 and 2 follows the first major problem of PCA-clustering: according to the Lemmas, the kernel of the projection operator is non-empty when $d<D$ (Lemma 1), and any perturbations inside of the kernel do not have any effect on the loss (Lemma 2). Hence, loss function in (8-11) is a semi-norm, not penalizing errors in the directions that are orthogonal, or close to orthogonal with respect to the manifolds $\mathcal{T}_{k}$. As will be demonstrated on examples below (see, e.g., Fig. 2C and the left panel of Fig. 3C), this can lead to the PCA-clustering minimizers $\mathcal{T}_{k}$ that become orthogonal, or close to orthogonal, with respect to the actual nonlinear low-dimensional manifold - hereby ignoring the actual manifold data, and achieving a small value of $\mathcal{L}$ at the same time. The second problem of PCA-clustering is induced by the exact $0/1$ minimizer in the step (ii) described above: it confines the resulting manifold reconstruction $X^{\textrm{rec}}(:,t)$ in (2) to piece-wise linear functions only, not allowing to achieve smooth convex interpolations of nonlinear manifolds that should be possible with the original optimization problem formulation (4-7). But, this original formulation (4-7) does not allow obtaining the cheaply-computable analytic linearly-scalable solutions (i)-(ii) that are used in the PCA-clustering algorithm. And, solving the original problem (4-7) - being highly-nonlinear, subject to multivariate and polynomial constraints - with standard numerical tools from optimization theory would require very expensive numerical algorithms that have the exponential worst-case scaling [19].
@@ 2.2 Entropy-Optimal Manifold Clustering (EOMC)
To mitigate these two central problems of PCA-clustering, the two following modifications of the loss function $\mathcal{L}$ in (8-11) will be adopted:
  • Loss metrisation: to fix the problem identified by the Lemmas 1 and 2, we will introduce an additional scalar hyper-parameter $\alpha\geq 0$ and add an addtional term $\alpha\|X(:,t)-\mu_{k}\|_{2}^{2}$ to the loss of each of the $t$ instances for all $k=1,\dots,K$. As will be proven below, this modification upgrades the semi-norm loss of (8-11) to the $\alpha$-weighted norm, fixing the problem with the data points in the directions that are close to orthogonal to $\mathcal{T}_{k}$.
  • Entropy regularization: to avoid the problem with a piece-wise linearity restriction - and to allow for the $\gamma$-solutions to be not only $0/1$ (and, to achieve this without returning back to the original formulation (4-7) that would result in exponentialy-scaling worst case numerics) - we can use the fact that the columns of $\gamma$ can be interpreted as $t$-dependent $K$-dimensional probability measures. We can add the normalized negative Shannon entropy regularization terms $\beta\gamma(k,t)\textrm{log}_{K}\left(\gamma(k,t)\right)$ (with a scalar hyper-parameter $\beta\geq 0$) to the loss, to optimally-adjust the entropies of these probability measures during learning [^2] Base $K$ of the logarithm is used to normilize the entropy values to the interval $\left[0,1\right]$. This modification is motivated by the entropic learning methods that allow training the entropy-optimal probability measures by tuning this hyper-parameter $\beta$ in the loss. The $0/1$ $\gamma$-solution of the PCA-clustering is a minimum entropy measure, and is attained when setting $\beta=0$ [20, 21, 22, 23, 24].
@@
$\left\{\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}^{\textrm{EOMC}},$ (17)
$\mathcal{L}^{\textrm{EOMC}}$ $=$ $\frac{1}{T}\sum_{k,t=1}^{K,T}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right],$ (18)
@@
Please note that the PCA-clustering [16, 17, 18] is a special case of (17-20), attained when setting $\alpha=0$ and $\beta=0$. Standard linear PCA is also a special case of (17-20), when selecting $K=1,\alpha=0,\beta=0$.
@@
$x^{\dagger}A^{\alpha}x$ $=$ $x^{\dagger}\left((1+\alpha)I_{D}-P\right)x,$ (25)
$x^{\dagger}A^{\alpha}x$ $=$ $(1+\alpha)x^{\dagger}I_{D}x-x^{\dagger}Px,$ (26)
$x^{\dagger}A^{\alpha}x$ $=$ $(1+\alpha)\|x\|^{2}-\|Px\|^{2}.$ (27)
@@
$x^{\dagger}A^{\alpha}x$ $=$ $\|x\|^{2}+\alpha\|x\|^{2}-\|Px\|^{2},$ (28)
$x^{\dagger}A^{\alpha}x$ $=$ $(\|x\|^{2}-\|Px\|^{2})+\alpha\|x\|^{2}.$ (29)
@@
Finally, as proven in the Theorem 1 below, optimization problem (17-20) can be solved through a sequence of analytically-solvable steps - and without exceeding the computational cost of the original PCA-clustering algorithm for (8-11), i.e., with the overall leading iteration cost scaling of $\mathcal{O}(T)$.
@@
Let $X\in\mathbb{R}^{D,T}$, and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$ are fixed. Then, the EOMC problem (17-20) has the following properties:
@@
    Step 1 of EOMC-algorithm, analytic solutions for $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$: if $\sum_{t=1}^{T}\gamma(k,t)>0$, then for a fixed $\gamma$ that satisfies the constraints (20), the solution of (17-20) is provided by:
@@
    where $\underset{d}{\overline{\mathrm{eigvec}}}\left(A\right)$ denotes an operation of computing $d$ dominant eigenvectors of A (i.e., the eigenvectors that correspond to the $d$ largest eigenvalues of the symmetric positive-semidefinite operator $A$) and putting them in the $D\times d$ orthogonal matrix column-wise. Operator $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ is defined as $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)=\frac{\sum_{t=1}^{T}\gamma(k,t)\left(X(:,t)-\mu_{k}^{*}\right)\left(X(:,t)-\mu_{k}^{*}\right)^{\dagger}}{\sum_{t=1}^{T}\gamma(k,t)}$. Cost of (30-31) computation scales as $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$.
@@
    Step 2 of EOMC-algorithm, analytic solutions for $\gamma(k,t)$: let $g(t)=\left(g_{1}(t),\dots,g_{K}(t)\right)$, where $g_{k}(t)=\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)$, and let $g_{\min}(t)$ be the infinum of $g(t)$ for a fixed $t$. Then, for fixed $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ and when $\beta>0$, the solution of (17-20) is given by:
@@
    If $\beta=0$, then the solution of (17-20) is given by $\gamma^{*}(k^{*},t)=1$ for $k^{*}=\arg\min_{k}\left(g_{k}(t)-g_{\min}(t)\right)$, and $\gamma^{*}(k,t)=0$ for all $k\neq k^{*}$. Cost of computing $\gamma^{*}$ with (30-31) scales as $\mathcal{O}\left(TK\left(D+1\right)\right)$.
@@
    Monotonicity and convergence of EOMC-algorithm: starting with a randomly-generated $\gamma$ that satisfies constraints in (17-20), and iteratively repeating the above Step 1 and Step 2, results in monotonic decrease of the function $\mathcal{L}^{\textrm{EOMC}}$, converging to a local optimum of (17-20). The overall computational iteration cost of EOMC-algorithm scales as $\mathcal{O}\left(TK\left(D+1\right)^{2}+KdD^{2}\right)$
@@
    For fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$, for every $\mu_{k}$, $k=1,\dots,K$, according to the Lemma1 the problem (17-20) is a strictly-convex unconstrained problem with a unique minimum
    $\sum_{t=1}^{T}\gamma(t,k)\left(X(:,t)-\mu_{k}^{*}\right)=0,$ (34)
    that after re-arranging of terms provides (30) if $\sum_{t=1}^{T}\gamma(k,t)>0$. If $\sum_{t=1}^{T}\gamma(k,t)=0$, then any value of $\mu_{k}^{*}$ is a solution. Next, we fix these $\mu_{k}^{*}$, for all $k=1,\dots,K$, and consider the solution of (17-20) with respect to $\mathcal{T}_{k}$, for fixed $X\in\mathbb{R}^{T,D}$, $\gamma$ and the hyper-parameters $K\geq 1$, $d\leq K$, $\alpha>0$ and $\beta\geq 0$. Defining the $d\times d$ matrix of Lagrange multipliers $\Lambda$, we can rewrite (17-20) in the Euler-Lagrange form
@@
    $\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}$ $=$ $I_{d},$
    for all $k=1,\dots,K$. Substituting (36) into (35), taking trace and deploying algebraic transformations that make use of the trace operation properties, we obtain that $\Lambda$ in (36) is a diagonal matrix of dominant eigenvalues of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (trace of $\Lambda$ should maximize the trace of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ after projection on the orthogonal subspace $\mathcal{T}_{k}$). Hence, $\mathcal{T}_{k}$ corresponds column-wise to the $d$ dominant eigenvectors of $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ and $\mathcal{T}_{k}^{*}=\underset{d}{\overline{\mathrm{eigvec}}}\left(\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)\right)$. Summing up the cost of computing the weighted means and weighted covariances for each cluster with the cost of computing the $d$-dominant eigenvectors of symmetric positive-semidefinite matrix $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ (that is less then a cost of full diagonalization, and allows applying iterative Krylov methods with cost scaling of $\mathcal{O}\left(KdD^{2}\right)$), we obtain the cost scaling of $\mathcal{O}\left(TKD\left(D+1\right)+KdD^{2}\right)$ for Step 1 .
@@
    These properties follow from applying Lemma 2.1 and Lemma 2.6 in [24] to (17-20). The cost of computing Step 2 consists of $\mathcal{O}\left(TKD\right)$ (computing $K$ elements of vector $g(t)$ for all $t$), and adding the costs of $\mathcal{O}\left(TK\right)$ computations of (32) for every $t$.
@@
    As follows from the above proofs for items (1.) and (2.), each of the Step 1 and Step 2 leads to a monotonic decrease of the function value $\mathcal{L}^{\textrm{EOMC}}$ in (17-20). It is straightforward to validate that this sequence is bounded from below with $\mathcal{L}^{\textrm{EOMC}}\geq-\beta$. Hence, since iterative repetition of these steps also generates a monotonically-decreasing sequence - and it is bounded from below with $\beta$ in $\mathcal{R}^{1}$ - this monotonically-decreasing sequence is converging to some finite $\mathcal{L}^{\textrm{EOMC}}$ if $\beta<\infty$, s.t. $\|\mathcal{L}^{\textrm{EOMC}}\|<\infty$.